# REG4Rec: Reasoning-Enhanced Generative Model for Large-Scale Recommendation Systems

### v3

@@ 3.3.1. MPQ: Moe-based Parallel Quantization Codebooks
(2) $\mathcal{L}_{\text{recon}}=\left\|\mathbf{v}_{i}-\text{decoder}(\mathbf{q}_{i})\right\|^{2}_{2},$
@@
(3) $\mathcal{L}_{\text{orth}}$ $=\left\|(\bar{W})^{\top}\bar{W}-\mathbf{I}\right\|_{F}^{2},$
where $\mathbf{I}$ is the identity matrix and $\|\cdot\|_{F}^{2}$ denotes the squared Frobenius norm (Golub and Van Loan, 2013). The overall objective combines these terms:
(4) $\mathcal{L}=\mathcal{L}_{\text{recon}}+\alpha\mathcal{L}_{\text{orth}},$
@@
3.3.2. CRSS: Confidence-based Reasoning Step Selection
@@
(5) $c_{j}=\arg\max_{C^{k}\in\mathcal{C_{\text{val}}}}(\arg\max_{c\in C^{k}}\mathbf{S}(c)).$
@@
CRSS leverages multiple parallel codebooks in MPQ and uses confidence scores as a proxy for reliability, balancing accuracy and exploration during reasoning. This approach is well-suited to GR tasks in recommendation systems, where user behavior is highly variable and non-stationary (Zhang et al., 2023), and adaptive decoding is required to capture dynamic intents.
3.4. Training for Reliable Reasoning Process
@@ 3.4.1. Pretrain task
(6) $\mathcal{L}_{\text{token}}=-\sum_{c=1}^{M}\log p(c^{T+1}_{i}\mid c^{T+1}_{<i},\mathcal{S}_{u}),$
@@
(7) $\mathcal{L}_{\text{cate}}=-\log\tilde{p}(g_{i}^{\text{gr}}\mid c^{T+1}_{<i},\mathcal{S}_{u}),$
@@
3.4.2. PARS: Preference Alignment for Reasoning
@@
(8) $A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},...,r_{G}\})}{\text{std}(\{r_{1},r_{2},...,r_{G}\})},$
where $r_{i}$ is the cumulative reward of path $y_{i}$. Following GRPO algorithm (Shao et al., 2024), which employs policy optimization with Kullback-Leibler (KL) divergence (Kullback and Leibler, 1951) for stability, we define the objective function as follows:
$J(\theta)=\mathbb{E}_{u\sim\mathcal{S}_{u},\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(r_{i}A_{i},\ \text{clip}\left(r_{i},\ 1-\epsilon,\ 1+\epsilon\right)A_{i}\right)\right]$
where $r_{i}=\frac{\pi_{\theta}(y_{i}|u)}{\pi_{\theta_{\text{old}}}(y_{i}|u)}$, $\pi_{\theta_{\text{old}}}$ denotes the reference policy, $D_{\text{KL}}(\cdot||\cdot)$ represents the KL divergence measuring the discrepancy between the current policy and the reference policy.
@@
(10) $\gamma_{\text{step}}=\frac{1}{M}\sum_{i=1}^{M}\mathbb{I}(c_{i}^{\text{gr}}\in\mathcal{C}^{i}),$
@@
(11) $\gamma_{\text{cate}}=\frac{1}{M}\sum_{i=1}^{M}\mathbb{I}(g_{i}^{\text{gr}}=g_{i}^{\text{gt}}),$
@@
Step-Consistency Reward ($\gamma_{\text{js}}$): To promote semantic coherence across consecutive steps, we calculate the Jensen-Shannon (JS) divergence (Lin, 2002) between the predicted category distributions $\mathcal{G}$ at steps $i$ and $i+1$:
(12) $JS(\tilde{p}_{i}\|\tilde{p}_{i+1})=\frac{1}{2}D_{\text{KL}}(\tilde{p}_{i}\|m)+\frac{1}{2}D_{\text{KL}}(\tilde{p}_{i+1}\|m),$
@@
(13) $\gamma_{\text{js}}=1-\frac{1}{M-1}\sum_{i=1}^{M-1}JS(\tilde{p}_{i}\|\tilde{p}_{i+1}).$
@@
Global Path Reward ($\gamma_{\text{path}}$): We evaluate path robustness with a progressive retrieval scheme that allows up to $d$ tokens to be incorrect, testing the model’s ability to maintain retrieval performance under partial errors. For a reasoning path of length $M$, we build a cumulative sum of codebook representation $z$ using only the MdM-d correct tokens, $z=\sum_{m=1}^{M-d}z^{m}$, and perform top-NN retrieval from the item pool to check whether the target item is still retrieved. The reward under $d$ errors is:
(14) $\gamma_{\text{path}}=\begin{cases}1&\mathbb{I}(i_{T+1}\in\text{Top-}N(z)),\\ 0&\text{otherwise}.\end{cases}$
In Eq. (14), the indicator function $\mathbb{I}(\cdot)$ return 1 if the target item is ranked within the top-NN retrieved candidates and 0 otherwise.
@@
3.4.3. MSRA: Multi-Step Reward Augmentation Strategy
Traditional reward designs align the model with a single next item $i_{T+1}$. However, this strict one-step alignment often fails to capture the complexities of real-world sequential behaviors (Deng et al., 2025b). In real-world recommendation system, user intent emerges over short horizons and is noisy and dynamic. To address these challenges, we propose a MSRA Strategy that expands valid targets to future interactions within a short time window.
@@ 3.4.3. MSRA: Multi-Step Reward Augmentation Strategy
(15) $\gamma_{\text{step}}^{\text{msra}}=\frac{\sum_{j=1}^{h}w^{j-1}\gamma_{\text{step}}^{(j)}}{\sum_{j=1}^{h}w^{j-1}},$
@@
Additionally, We also extend the global path reward. Rather than requiring exact retrieval of the single target $i_{t+1}$, the augmented path reward is positive if the top-NN results include any of the $h$ candidate items $\{i_{t+1},\dots,i_{t+h}\}$. This relaxation encourages the model to accommodate broader user intents. Formally, let $W_{t}=\{i_{t+1},\dots,i_{t+h}\}$ represent the set of $h$ candidate items, and let $\text{Top-}N(z)$ be the top-NN retrieved items using the cumulative representation $z$ of the generated path:
@@
(16) $\gamma_{\text{path}}^{\text{msra}}=\frac{\sum_{j=1}^{h}\left(w^{j-1}\mathbb{I}(i_{t+j}\in\text{Top-}N(z))\right)}{\sum_{j=1}^{h}w^{j-1}},$
@@
where $\mathbb{I}(\cdot)$ denotes an indicator function that returns 1 if any item in the candidate set is included in the top-NN results and 0 otherwise. By applying time-decayed weighting, MSRA prioritizes recent interactions while preserving sensitivity to long-term preferences, improving the robustness and diversity of the reasoning paths.
3.4.4. Efficient Training Strategy
While post-training improves REG4Rec’s ability to select reliable reasoning paths, it increases offline training cost. Conventional pipelines typically use a uniform, fixed numerical format for all layers—fp32 (full-precision floating point), bf16 (half-precision floating point), or fp8 (an 8-bit floating‑point format)—thus ignoring layer-wise heterogeneity and temporal sensitivity (Rouhani et al., 2023; Li et al., 2024; Wang et al., 2018). To cope with resource constraints in large-scale recommender systems, we propose Layer-Adaptive Dynamic Quantization Controller (LADQ), which optimizes the quantization precisions for each layer to accelerate training.
Our approach is motivated by a Hessian-based sensitivity analysis (Dong et al., 2020), which revealed two key observations: (i) certain layers are highly quantization-sensitive, where low-precision quantization drastically degrades model accuracy; and (ii) this sensitivity is not static but evolves dynamically as training progresses. Informed by these insights, LADQ operates adaptively: during training it periodically re-evaluates each layer’s sensitivity and then chooses the best precision among {fp32, bf16, fp8} by jointly considering both sensitivity and the layer’s contribution to training latency. This layer-wise, adaptive strategy significantly accelerates the post-training phase while minimizing the trade-off in performance.
3.5. Inference with Reflection

### v4

@@ IV-1 MPQ: Moe-based Parallel Quantization Codebooks
$\mathcal{L}_{\text{recon}}=\left\|\mathbf{v}_{i}-\text{decoder}(\mathbf{q}_{i})\right\|^{2}_{2},$ (2)
@@
$\mathcal{L}_{\text{orth}}$ $=\left\|(\bar{W})^{\top}\bar{W}-\mathbf{I}\right\|_{F}^{2},$ (3)
where $\mathbf{I}$ is the identity matrix and $\|\cdot\|_{F}^{2}$ denotes the squared Frobenius norm [42]. The overall objective combines these terms:
$\mathcal{L}=\mathcal{L}_{\text{recon}}+\alpha\mathcal{L}_{\text{orth}},$ (4)
@@
IV-2 CRSS: Confidence-based Reasoning Step Selection
@@
$c_{j}=\arg\max_{C^{k}\in\mathcal{C_{\text{val}}}}(\arg\max_{c\in C^{k}}\mathbf{S}(c)).$ (5)
@@
CRSS leverages multiple parallel codebooks in MPQ and uses confidence scores as a proxy for reliability, balancing accuracy and exploration during reasoning. This approach is well-suited to GR tasks in recommendation systems, where user behavior is highly variable and non-stationary [32], and adaptive decoding is required to capture dynamic intents.
Algorithm 1 Pre-training for Generative Recommendation
1:User-item interaction data $\mathcal{D}=\{(u,\mathcal{S}_{u})\}$, item feature vectors $\{\mathbf{v}_{i}\}_{i\in\mathcal{I}}$, number of parallel experts $M$, codebook size $D$, number of decoding steps $M$.
2:Parallel autoencoder experts $\{E_{r}\}_{r=1}^{M}$, codebooks $\{\mathcal{C}^{r}=\{z^{r}_{1},\dots,z^{r}_{D}\}\}_{r=1}^{M}$, projection matrices $\{W_{r}\}_{r=1}^{M}$, shared decoder, Transformer encoder-decoder model $\pi_{\theta}$, category classifier $f_{c}$, pre-trained feature extractors $\{\phi_{t},\phi_{v},\phi_{b}\}$.
3:
4:while not converged do
5:  for all item $i\in\mathcal{I}$ do
6:    $\mathbf{v}_{i}\leftarrow[\phi_{t}(i),\phi_{v}(i),\phi_{b}(i)]$
7:    for $r=1$ to $M$ do
8:     $e^{r}_{i}\leftarrow E_{r}(\mathbf{v}_{i})$
9:     $c^{i}_{r}\leftarrow\arg\min_{j}\|e^{r}_{i}-z^{r}_{j}\|_{2}$
10:     $z^{r}_{i}\leftarrow z^{r}_{c^{i}_{r}}$
11:    end for
12:    Compute routing weights: $h^{r}(\mathbf{v}_{i})\leftarrow\text{MoE-gate}(\mathbf{v}_{i})$
13:    $\mathbf{q}_{i}\leftarrow\sum_{r=1}^{M}h^{r}(\mathbf{v}_{i})\cdot z^{r}_{i}$
14:    $\mathcal{L}_{\text{recon}}\leftarrow\left\|\mathbf{v}_{i}-\text{decoder}(\mathbf{q}_{i})\right\|_{2}^{2}$
15:    Normalize matrices: $\bar{W}\leftarrow\text{L2-norm}([W_{1},\dots,W_{M}])$
16:    $\mathcal{L}_{\text{orth}}\leftarrow\left\|\bar{W}^{\top}\bar{W}-\mathbf{I}\right\|_{F}^{2}$
17:    $\mathcal{L}\leftarrow\mathcal{L}_{\text{recon}}+\alpha\cdot\mathcal{L}_{\text{orth}}$
18:    Update $\{E_{r},\mathcal{C}^{r},W_{r},\pi_{\theta}\}$ via gradient descent on $\mathcal{L}$
19:  end for
20:end while
21:
22:while not converged do
23:  Sample $(u,\mathcal{S}_{u})\sim\mathcal{D}$
24:  Encode history: $H\leftarrow\text{Encoder}(\mathcal{S}_{u})$
25:  Decode token sequence $y=[c_{1},\dots,c_{M}]$
26:  Compute token loss: $\mathcal{L}_{\text{token}}\leftarrow-\sum_{i=1}^{M}\log p(c_{i}\mid c_{<i},H)$
27:  Predict categories: $g_{i}^{\text{gr}}\leftarrow f_{c}(c_{<i},H),\ \forall i$
28:  Compute category loss: $\mathcal{L}_{\text{cate}}\leftarrow-\log\tilde{p}(g_{i}^{\text{gr}}\mid c_{<i},H)$
29:  Update $\pi_{\theta}$ via gradient descent on $\mathcal{L}_{\text{token}}+\lambda\mathcal{L}_{\text{cate}}$
30:end while
V Training for Reliable Reasoning Process
@@ V-1 Pretrain task
$\mathcal{L}_{\text{token}}=-\sum_{c=1}^{M}\log p(c^{T+1}_{i}\mid c^{T+1}_{<i},\mathcal{S}_{u}),$ (6)
@@
$\mathcal{L}_{\text{cate}}=-\log\tilde{p}(g_{i}^{\text{gr}}\mid c^{T+1}_{<i},\mathcal{S}_{u}),$ (7)
@@
V-2 PARS: Preference Alignment for Reasoning
@@
$A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},...,r_{G}\})}{\text{std}(\{r_{1},r_{2},...,r_{G}\})},$ (8)
where $r_{i}$ is the cumulative reward of path $y_{i}$. Following GRPO algorithm [22], which employs policy optimization with Kullback-Leibler (KL) divergence [28] for stability, we define the objective function as follows:
$\begin{split}J(\theta)&=\mathbb{E}_{s_{u}\sim D,\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|s_{u})}\\ &\quad\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\frac{\pi_{\theta}(y_{i}|s_{u})}{\pi_{\theta_{\text{old}}}(y_{i}|s_{u})}A_{i},\right.\right.\\ &\quad\quad\left.\left.\text{clip}\left(\frac{\pi_{\theta}(y_{i}|s_{u})}{\pi_{\theta_{\text{old}}}(y_{i}|s_{u})},1-\epsilon,1+\epsilon\right)A_{i}\right)\right.\\ &\quad\quad\left.-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\theta_{\text{ref}}})\right],\end{split}$ (9)
where $\pi_{\theta_{\text{old}}}$ denotes the reference policy, $D_{\text{KL}}(\cdot||\cdot)$ represents the KL divergence measuring the discrepancy between the current policy and the reference policy.
@@
$\gamma_{\text{step}}=\frac{1}{M}\sum_{i=1}^{M}\mathbb{I}(c_{i}^{\text{gr}}\in\mathcal{C}^{i}),$ (10)
@@
$\gamma_{\text{cate}}=\frac{1}{M}\sum_{i=1}^{M}\mathbb{I}(g_{i}^{\text{gr}}=g_{i}^{\text{gt}}),$ (11)
@@
Step-Consistency Reward ($\gamma_{\text{js}}$): To promote semantic coherence across consecutive steps, we calculate the Jensen-Shannon (JS) divergence [35] between the predicted category distributions $\mathcal{G}$ at steps $i$ and $i+1$:
$JS(\tilde{p}_{i}\|\tilde{p}_{i+1})=\frac{1}{2}D_{\text{KL}}(\tilde{p}_{i}\|m)+\frac{1}{2}D_{\text{KL}}(\tilde{p}_{i+1}\|m),$ (12)
@@
$\gamma_{\text{js}}=1-\frac{1}{M-1}\sum_{i=1}^{M-1}JS(\tilde{p}_{i}\|\tilde{p}_{i+1}).$ (13)
@@
Global Path Reward ($\gamma_{\text{path}}$): We evaluate path robustness with a progressive retrieval scheme that allows up to $d$ tokens to be incorrect, testing the model’s ability to maintain retrieval performance under partial errors. For a reasoning path of length $M$, we build a cumulative sum of codebook representation $z$ using only the MdM-d correct tokens, $z=\sum_{m=1}^{M-d}z^{m}$, and perform top-$N$ retrieval from the item pool to check whether the target item is still retrieved. The reward under $d$ errors is:
$\gamma_{\text{path}}=\begin{cases}1&\mathbb{I}(i_{T+1}\in\text{Top-}N(z)),\\ 0&\text{otherwise}.\end{cases}$ (14)
In Eq. (14), the indicator function $\mathbb{I}(\cdot)$ return 1 if the target item is ranked within the top-$N$ retrieved candidates and 0 otherwise.
@@
V-3 MSRA: Multi-Step Reward Augmentation Strategy
Traditional reward designs align the model with a single next item $i_{T+1}$. However, this strict one-step alignment often fails to capture the complexities of real-world sequential behaviors [5]. In real-world recommendation system, user intent emerges over short horizons and is noisy and dynamic. To address these challenges, we propose a MSRA Strategy that expands valid targets to future interactions within a short time window.
@@ V-3 MSRA: Multi-Step Reward Augmentation Strategy
$\gamma_{\text{step}}^{\text{msra}}=\frac{\sum_{j=1}^{h}w^{j-1}\gamma_{\text{step}}^{(j)}}{\sum_{j=1}^{h}w^{j-1}},$ (15)
@@
Additionally, We also extend the global path reward. Rather than requiring exact retrieval of the single target $i_{t+1}$, the augmented path reward is positive if the top-$N$ results include any of the $h$ candidate items $\{i_{t+1},\dots,i_{t+h}\}$. This relaxation encourages the model to accommodate broader user intents. Formally, let $W_{t}=\{i_{t+1},\dots,i_{t+h}\}$ represent the set of $h$ candidate items, and let $\text{Top-}N(z)$ be the top-$N$ retrieved items using the cumulative representation $z$ of the generated path:
@@
$\gamma_{\text{path}}^{\text{msra}}=\frac{\sum_{j=1}^{h}\left(w^{j-1}\mathbb{I}(i_{t+j}\in\text{Top-}N(z))\right)}{\sum_{j=1}^{h}w^{j-1}},$ (16)
@@
where $\mathbb{I}(\cdot)$ denotes an indicator function that returns 1 if any item in the candidate set is included in the top-$N$ results and 0 otherwise. By applying time-decayed weighting, MSRA prioritizes recent interactions while preserving sensitivity to long-term preferences, improving the robustness and diversity of the reasoning paths.
Algorithm 2 Post-training for Generative Reasoning
1:User-item sequences $\mathcal{D}=\{(u,\mathcal{S}_{u})\}$, MPQ codebooks $\{\mathcal{C}^{r}\}$, decoding steps $M$, rollout number $G$, future window $h$, time-decay factor $w\in(0,1)$, clip range $\epsilon$, KL weight $\beta$
2:Policy $\pi_{\theta}$, old policy $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$, reference policy $\pi_{\theta_{\text{ref}}}\leftarrow\pi_{\theta}$, category classifier $f_{c}$
3:while not converged do
4:  Sample $(u,\mathcal{S}_{u})\sim\mathcal{D}$, encode history $H\leftarrow\text{Encoder}(\mathcal{S}_{u})$
5:  Sample $G$ rollouts: $\tau=\{y_{1},\dots,y_{G}\}\sim\pi_{\theta_{\text{old}}}(\cdot\mid H)$
6:  for all $y_{i}=[c_{1},\dots,c_{M}]\in\tau$ do
7:   Predict categories: $g_{k}^{\text{gr}}\leftarrow f_{c}(c_{<k},H),\ \forall k$
8:   Candidate items: $W=\{i_{t+1},\dots,i_{t+h}\}$
9:   $Z\leftarrow\sum_{j=1}^{h}w^{j-1}$, initialize $\gamma^{\text{msra}}\leftarrow 0$
10:   for $j=1$ to $h$ do
11:     $\gamma_{\text{step}}^{(j)}\leftarrow\frac{1}{M}\sum_{k=1}^{M}\mathbb{I}(c_{k}\in\mathcal{C}^{k}(i_{t+j}))$
12:     $\gamma_{\text{cate}}^{(j)}\leftarrow\frac{1}{M}\sum_{k=1}^{M}\mathbb{I}(g_{k}^{\text{gr}}=g^{\text{gt}}(i_{t+j}))$
13:     $z\leftarrow\sum_{k=1}^{M}z_{k}$, $\mathbb{I}_{\text{path}}^{(j)}\leftarrow\mathbb{I}(i_{t+j}\in\text{Top-}N(z))$
14:     $\gamma^{\text{msra}}\mathrel{+}=w^{j-1}\cdot\left(\gamma_{\text{step}}^{(j)}+\gamma_{\text{cate}}^{(j)}+\mathbb{I}_{\text{path}}^{(j)}\right)$
15:   end for
16:   Normalize MSRA rewards: $\gamma^{\text{msra}}\leftarrow\gamma^{\text{msra}}/Z$
17:   $\gamma_{\text{js}}\leftarrow 1-\frac{1}{M-1}\sum_{k=1}^{M-1}JS(\tilde{p}_{k}\|\tilde{p}_{k+1})$
18:   Total reward: $r_{i}\leftarrow\gamma^{\text{msra}}+\gamma_{\text{js}}$
19:  end for
20:  Normalize advantages: $A_{i}\leftarrow(r_{i}-\mu_{r})/\sigma_{r}$
21:  Compute ratio: $\rho_{i}\leftarrow\pi_{\theta}(y_{i}\mid H)/\pi_{\theta_{\text{old}}}(y_{i}\mid H)$
22:  Update $\pi_{\theta}$ by gradient ascent on via Equation 9, then $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
23:end while
VI Deployment of REG4Rec
VI-A Efficient Training Strategy
While post-training significantly improves model performance, it introduces substantial offline training costs, a critical concern in large-scale recommender systems. A common acceleration strategy is to reduce numerical precision. However, conventional pipelines typically adopt a uniform, fixed format for all layers, such as fp32 (full-precision floating point), bf16 (half-precision floating point), or fp8 (an 8-bit floating‑point format), ignoring the inherent heterogeneity and temporal dynamics of different model components [38, 37, 40]. Our initial attempts to apply standard mixed-precision training (AMP) with BF16 resulted in noticeable accuracy degradation compared to the FP32 baseline. As shown in Figure 3, the loss trajectories for BF16 and FP8 plateaued prematurely in later training stages, while FP32 continued to converge.
Figure 3: Loss progression across precisions.
Through investigation, we identified two root causes for this performance drop: (i) Quantization Sensitivity: Activations and gradients in certain layers are highly sensitive to low precision, frequently underflowing to zero in BF16 and causing information loss. (ii) Insufficient Gradient Scaling: The standard gradient scaling in AMP was inadequate to compensate for the extremely small gradients in these sensitive layers. This motivated a more intelligent, fine-grained approach to quantization.
To balance computational efficiency and model accuracy, we propose the Layer-Adaptive Dynamic Quantization Controller (LADQ). The core principle of LADQ is to dynamically allocate higher precision to quantization-sensitive layers while assigning lower precision to robust, computationally-intensive layers. To implement this, we first needed a reliable method to measure layer-wise sensitivity during training. We explored several candidate metrics traditionally used for post-training quantization, including the Hessian matrix [36], the Fisher matrix, and the K-FAC matrix [44], which are conventionally applied for quantization sensitivity measurement during inference. We extended these approaches through sampling-based measurement during the training process.
Through repeated sampling during training, we derived two key findings that inform LADQ’s design:
  • Dynamic Distribution: The set of quantization sensitive layers is not static but shifts across different training phases.
  • Hybrid Sensitivity: Layer-wise sensitivity varies significantly. For example, as illustrated in Figure 4, within our model’s transformer architecture, module 6 (the attention mechanism in the second layer) exhibited pronounced sensitivity, particularly during late-stage training.
Figure 4: K-FAC matrix trace sampling during training.
To validate sensitivity metrics, we manually tested quantization configurations. Results confirmed that preserving high precision for sensitive modules (e.g., module 6) minimized loss degradation. Among Hessian, Fisher, and K-FAC, K-FAC proved optimal due to its superior accuracy-computation tradeoff.
Based on these insights, we designed the LADQ controller to automate precision allocation. This controller operates periodically throughout the training process, executing a three-step cycle:
  • (i). Computes per-layer sensitivity.
  • (ii). Calculates a reward score for each layer, which jointly considers sensitivity and contribution to overall training latency.
  • (iii). Dynamically assigns the optimal precision from the set {fp32, bf16, fp8} to each layer, prioritizing higher precision for high-reward (sensitive but critical) layers.
By adapting precision at a layer-wise level based on dynamic sensitivity, LADQ successfully accelerates the training phase while maintaining loss convergence on par with the full-precision baseline, providing a practical solution for resource-constrained environments.
VI-B Inference with Reflection
Inspired by the consistency-oriented rewards in our GRPO framework, we propose the Consistency Oriented Self‑Reflection Pruning Strategy (CORP). This approach refines the standard beam search by incorporating a measure of reasoning consistency into the path selection process, thereby steering the generation towards more logically coherent sequences and mitigating the introduction of noise.
The core mechanism is as follows. During beam search, for each candidate path being extended, we do not only consider the probability of the next token. At every $s$-th step, the category predictor produces a probability distribution $\tilde{p}_{i}$ over item categories, serving as a semantic signal for consistency checking. This distribution acts as a snapshot of the model’s understanding or reasoning focus at this step.